solve the system using elimination: the problem is
6x+3y=-12 and 6x+2y= -4

Answer by checkley77(12844)

6x+3y=-12
-(6x+2y= -4) subtract.
-----------------
y=-8 ans.
6x+3*-8=-12
6x-24=-12
6x=-12+24
6x=12
x=12/6
x=2 ans.
Proof:
6*2+2*-8=-4
12-16=-4
-4=-4